tet 3. Let (1 3 3 A= 0 1 3 . f(t) = 0-3 1/ Find the general solution to i = AT + f. -3te2t \(t+11)e24
Answer #1

teet ا ل ا اعه * (t+11 Jet oo is Az 11 3 31 f(t) = \-ztet lo-3 il find the general som to š= Art F Solm for homogeneous systexp = (åt +B) ezt . X = 20²4 (åttb) + ezt a x = Art F 2e²t (@t+b) + 4 3 = Alat to geet + Ý. o = (A-2) (at +53e2t + ? - eztã. 0